How much work is required to stop an electron (m=9.11×10−31kg) which is moving with a speed of 1.40×106 m/s ?

Express your answer to three significant figures and include the appropriate units.
Answer #1

Given,

v = 1.4 x 10^6 m/s

the amount of energy required to stop the electron is equal to the KE.

W = KE = 1/2 m v^2

W = 1/2 x 9.11 x 10^-31 x (1.4 x 10^6)^2 = 8.93 x 10^-19 J

Hence, W = 8.93 x 10^-19 J